Compute the determinant of :
∣∣∣−2−314∣∣∣
1
Expert's answer
2014-04-10T05:45:41-0400
Using triangle rule we get
| -2 -3 |
| 1 4 | = -2*4-(-3)*1 = -8+3 = -5
Answer: -5.